OpenStudy (anonymous):

divide x^3/4 / x^1/6

OpenStudy (anonymous):

\[x ^{3/4}/x ^{1/6}\]

OpenStudy (anonymous):

Just divide m^n/m^q = m^(n-q)

OpenStudy (anonymous):

m as the nominator and n / q as the denominators?

OpenStudy (anonymous):

\[\huge \frac{m^p}{m^q}=m^{p-q}\]

OpenStudy (anonymous):

So, \[\huge \frac{x^{3/4}}{x^{1/6}}=x^{(\frac{3}{4}-\frac{1}{6})}\]

OpenStudy (anonymous):

okay, I get it thanks

OpenStudy (anonymous):

All right

OpenStudy (anonymous):

Cheers